A committee of 5 people is to be chosen from a group of 6 men and 4 women. How many committees are
possible if it must consist of a majority of women?

Answer by mccravyedwin(407)

Case 1: The committee has 3 women and 2 men.  That's (4C3)(6C2)
Case 2: The committee has 4 women and 1 men.  That's (4C4)(6C1)

(4C3)(6C2) + (4C4)(6C1) = (4)(15) + (1)(6) = 60 + 6 = 66 possible committees.

Edwin